OpenStudy (anonymous):

Could someone show me how to solve this? (n!)/(n-3)!=5000

OpenStudy (anonymous):

\[\frac{n!}{(n-3)!}=\frac{n(n-1)(n-2)(n-3)!}{(n-3)!}=n(n-1)(n-2)\]

OpenStudy (anonymous):

Yes well I need my result to be a whole number..... Like I need to solve for n

OpenStudy (anonymous):

Could you show me how to get be the number result for n?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Does anybody know?? ?

OpenStudy (anonymous):

@wolf1728 @campbell_st

OpenStudy (wolf1728):

(n!)/(n-3)!=5000 can be rewritten (n * n-1) / (n-3)! = 5000 I'm working on it

OpenStudy (campbell_st):

well the way I see it... you can remove common factors from the numerator and denominator and are left with the equation n(n-1)(n -2) = 5000 or \[n^3 - 3n^2 + 2n -5000 = 0\] so its a process of solving for n or finding 3 consecutive numbers that multiply to 5000

OpenStudy (anonymous):

What do you mean @campbell_st ?

OpenStudy (campbell_st):

ok... n! = n x (n -1) x (n -2) x (n -3) x(n -4) ....... x 2 x 1 (n -3)! = (n -3) x (n -4) x (n -5) x.....2 x 1 so if you eliminate common factors from the numerator and denominator you get n x (x -1) x (n -2) = 5000

OpenStudy (anonymous):

So what would that result to? Like what whole number would n equal. That is what i am trying to solve

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

the problem is this doesn't give n as a integer...

OpenStudy (anonymous):

I don't know..lol I am rusty in this @campbell_st

OpenStudy (campbell_st):

it means n = 18.1193

OpenStudy (anonymous):

hmm okay

OpenStudy (campbell_st):

are you sure its not an inequation...?

OpenStudy (anonymous):

I give up lol

OpenStudy (campbell_st):

well its not something like largest n so that \[\frac{n!}{(n -3)!} < 5000\]

OpenStudy (wolf1728):

Yes, I also get a non-interger solution too n=18.119

OpenStudy (campbell_st):

tough question the way its written.... I was expecting n was a positive integer

OpenStudy (wolf1728):

BaileyC - are you absolutely sure of the way the question is written?

OpenStudy (anonymous):

u r left with n(n-1)(n-2)=5000 n[n^2-2n-n+2]=5000 n[n^2-3n-2]=5000